# Personalized Subgraph Federated Learning with Differentiable Auxiliary Projections

### v1

@@ Abstract
Federated learning (FL) on graph-structured data typically faces non-IID challenges, particularly in scenarios where each client holds a distinct subgraph sampled from a global graph. In this paper, we introduce Federated learning with Auxiliary projections (FedAux), a personalized subgraph FL framework that learns to align, compare, and aggregate heterogeneously distributed local models without sharing raw data or node embeddings. In FedAux, each client jointly trains (i) a local GNN and (ii) a learnable auxiliary projection vector (APV) that differentiably projects node embeddings onto a 1D space. A soft-sorting operation followed by a lightweight 1D convolution refines these embeddings in the ordered space, enabling the APV to effectively capture client-specific information. After local training, these APVs serve as compact signatures that the server uses to compute inter-client similarities and perform similarity-weighted parameter mixing, yielding personalized models while preserving cross-client knowledge transfer. Moreover, we provide rigorous theoretical analysis to establish the convergence and rationality of our design. Empirical evaluations across diverse graph benchmarks demonstrate that FedAux substantially outperforms existing baselines in both accuracy and personalization performance.
@@ 1 Introduction
Real-world data often manifests as relational structures, ranging from social interactions (Tan et al., 2023) and financial networks (Suzumura et al., 2019; Zhuo et al., 2024) to molecular graphs (Xie et al., 2021; Zhuo & Tan, 2022a), whose scale and privacy constraints increasingly require training to be carried out in a federated manner (He et al., 2021), whereby multiple clients collaboratively learn a Graph Neural Network (GNN) model without exchanging their raw data. However, applying federated learning to graph-structured data, such as social networks, faces severe challenges due to non-identically and independently distributed (non-IID) data across clients. For example, consider a federated learning scenario involving multiple regional social networking platforms, each representing a distinct subgraph of a global social network. Users within each region exhibit unique interaction patterns and distinct interests, resulting in significant heterogeneity in local graph structures and node attributes. This inherent diversity among subgraphs leads to substantial difficulties when attempting to aggregate local GNN models into a unified global model, as traditional FL algorithms (McMahan et al., 2017; Li et al., 2020) typically assume homogeneous data distributions across clients.
To tackle the non-IID challenges inherent in subgraph federated learning, personalized FL (Tan et al., 2022) has recently emerged as a promising paradigm, which aims to provide client-specific GNN models rather than enforcing a universal global solution. Existing personalized subgraph FL approaches commonly achieve personalization by clustering clients on the server side, necessitating a reliable measure of client similarity without direct access to client-side data. In this work, we impose even stricter privacy constraints: neither raw data nor embeddings are shared, and only model learnable parameters can be exchanged. Although the server could compare clients by directly measuring similarity between their parameter matrices uploaded, the high dimensionality of these matrices makes such metrics unreliable under the curse of dimensionality (Bellman, 1966). Recent improvements have proposed measuring similarity by comparing communication-level parameter gradients (Xie et al., 2021) or generating a common anchor graph on the server as a neutral testbed (Baek et al., 2023). Although these strategies mitigate some limitations, they remain largely heuristic and do not explicitly model the heterogeneity inherent in subgraph clients (See extended discussion in Appendix A).
Figure 1: The overall framework of FedAux.
@@
Contribution In this work, we propose FedAux, which employs differentiable auxiliary projections to effectively capture and exploit client-specific heterogeneity for subgraph FL. As illustrated on the left of Fig. 1, the server stores not only a global GNN but also a learnable auxiliary projection vector (APV) that accompanies the model parameters. At the start of the first communication round, the server broadcasts the global GNN and the current APV to all clients. Each client projects its node embeddings onto the APV, which is treated as a one-dimensional latent space. A differentiable soft-sorting operator then orders the projected embeddings by similarity, after which a simulated 1D convolution aggregates the sorted embedding sequence. The aggregated representations drive a supervised loss that simultaneously refines the local GNN and the APV, so that the optimized APV preserves the relational structure of the client subgraph. Upon completing local training, clients send their updated GNN weights and personalized APVs back to the server. Since the APV reveals only the latent space that best preserves local node relationships while concealing the exact position of every node in this space, it acts as a compact privacy-preserving summary of the client subgraph. Then the server computes similarity among the returned APVs to quantify inter-client affinity and yields client-specific aggregation weights. The server uses these weights to combine the incoming parameters, producing a personalized model for each client that respects both shared knowledge and local subgraph idiosyncrasies.
@@ 2 Problem Statement: Subgraph Federated Learning
In federated learning (FL), multiple clients collaboratively train a global model without exchanging their raw data. In the subgraph federated learning setting, each client holds a subgraph of a larger graph. Formally, let a graph $\mathcal{G}$ be partitioned (or sub-sampled) into $K$ subgraphs $\{G_{1},G_{2},\cdots,G_{K}\}$ as $K$ clients, where $G_{k}=(V_{k},E_{k},X_{k},Y_{k})$. Here, $V_{k}=\{v_{k,1},\cdots,v_{k,N_{k}}\}$ is the set of nodes in the $k$-th subgraph with $N_{k}=|V_{k}|$ nodes, $E_{k}$ the set of edges among those nodes, $X_{k}\in\mathbb{R}^{N_{k}\times d}$ the node features, and $Y_{k}$ the labels relevant to the learning task. In our FL scenario, each client $G_{i}$ has access only to its local data (i.e., its subgraph structure, node features, and labels), and there is no sharing of raw data or any node embeddings between clients.
A typical GNN $f_{\theta_{k}}(G_{k})$ parameterized by $\theta_{k}$ is employed to produce node embeddings and ultimately generate predictions on the client $G_{k}$. In a standard federated learning setting such as FedAvg (McMahan et al., 2017), one aims to solve the global objective: $\min_{\theta}\sum_{k=1}^{K}\alpha_{k}\mathcal{L}_{k}(\theta)$, subject to the privacy constraint that raw local data $G_{k}$ never leaves the client side. A common choice is to weight client $G_{k}$ by $\alpha_{k}=N_{k}/\bigl{(}\sum_{j=1}^{K}N_{j}\bigr{)}$ or simply $\alpha_{k}=1/K$. The iterative procedure proceeds as follows. First, the server initializes $\theta^{(0)}$. At each global communication round $t\in\{1,\cdots,T\}$, it sends $\theta^{(t-1)}$ to each client. $G_{k}$ then updates $\theta^{(t-1)}$ locally by taking a few stochastic gradient steps on $\mathcal{L}_{k}(\theta)$ to update the parameters $\theta_{k}\leftarrow\theta_{k}-\eta\nabla\mathcal{L}$, which produce $G_{k}$’s optimal local parameters $\theta_{k}^{(t)}$. After the $t$-th local training, all clients’ locally updated parameters $\{\theta_{1}^{(t)},\cdots\theta_{K}^{(t)}\}$ are sent back to the server, which aggregates them via a weighted average $\theta^{(t)}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}^{(t)}$. The newly aggregated global parameters $\theta^{(t)}$ are then broadcast back to each client for the next communication round. When the process converges or reaches a designated number of rounds, the final global parameters $\theta^{(T)}$ are taken as the parameters of the GNN model on the server.
@@ 3 Methodology
In this section, we introduce the proposed Subgraph Federated Learning with Auxiliary Projections (FedAux) framework, designed to address the heterogeneity across local subgraphs in federated learning. Fig. 1 illustrates an overview of FedAux. Our objective is twofold: 1) Each client locally encodes its subgraph into a one-dimensional space via a learnable auxiliary projection vector APV, and 2) the server then exploits these auxiliary vectors to realize personalized aggregation.
@@ 3.1 Client-Side Local Training
[…]
@@
which aligns node embeddings according to their coordinates in the $\bm{a}_{k}$-space. As shown in Fig. 2, during local training, the node embeddings and the structure of the $\bm{a}_{k}$-space are jointly optimized so that nodes with stronger relationships are positioned closer along this learned space (in $G_{k}$ comprising two triangles, nodes within the same triangle should be proximate in the $\bm{a}_{k}$-space). In other words, the objective of the local model is to adaptively reshape the APV so that the induced node sorting effectively captures the local data information.
[…]
@@
where $\mathbf{W}_{\tau}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ are learnable convolution kernels for offset $\tau$, and $b$ is a bias term. Convolving around $v_{\pi_{k}(i)}$ in Eq. 4 amounts to a proximity-based aggregation in the $\bm{a}_{k}$-space, where each embedding is updated by aggregating information from its neighbors along this learned 1D sorting. Hence, the quality of this sorting significantly impacts the aggregation effectiveness. Intuitively, nodes with stronger semantic relationships or similar labels should appear closer together in this learned sorting. We can formulate the learning objective to explicitly optimize the sorting induced by the APV $\bm{a}_{k}$, ensuring that the resultant sorting facilitates effective aggregation and improves downstream predictive accuracy. Consequently, we formulate the learning objective:
@@
$\left(\theta_{k}^{*},\bm{a}_{k}^{*},\Phi_{k}^{*}\right)=\underset{\theta_{k},a_{k},\Phi_{k}}{\operatorname{argmin}}\mathcal{L}\left(\operatorname{CLF}\left(\operatorname{Conv}1\mathrm{D}\left(\tilde{\mathbf{H}}_{k}^{(t-1)}\right)\right),Y_{k}\right),$ (5)
@@
where $\Phi_{k}$ denotes the full set of parameters for Conv1D and the subsequent classifier $\operatorname{CLF}$ that maps node embeddings to final logits. Through this objective, we explicitly encourage $\bm{a}_{k}$-space to yield an optimal node sorting, enabling the convolutional operation to effectively capture and leverage localized, label-informed relationships, thereby the optimized APV $\bm{a}^{(t)}_{k}=\bm{a}^{*}_{k}$ accurately preserves and encodes the local node relationships specific to each client.
@@
However, $\bm{a}_{k}$ does not directly participate in the loss defined in Eq. 5 in a way that enables standard backpropagation to update it. It is because the role of $\bm{a}_{k}$ is limited to generating similarity scores, which in turn determine the input order to the Conv1D layer. Thus, $\bm{a}_{k}$ only affects the network’s output by reordering embeddings, which is a purely indirect pathway that does not produce a gradient signal for $\bm{a}_{k}$ from the downstream loss.  Liu et al. (2021) attempted to mitigate this by multiplying each node embedding by its similarity score and then sorting. While this modification integrates $\bm{a}_{k}$ into the learning pipeline directly, the hard discrete sort persists, causing the gradient signal that could refine $\bm{a}_{k}$ to be still routed through a non-smooth transformation. Hence $\bm{a}_{k}$ still cannot be fully optimized to reorder embeddings based on loss feedback, leaving the core issue unresolved.
@@
To eliminate the hard-sorting bottleneck, we propose a continuous aggregation scheme over the $\bm{a}_{k}$-space. Rather than ranking or discretizing these similarity scores, for each node $v_{i}$, we define a continuous kernel $\kappa(s^{(t-1)}_{k,i},s^{(t-1)}_{k,j})$, which could be a simple Gaussian-like function $\mathcal{K}_{ij}=\kappa(s^{(t-1)}_{k,i},s^{(t-1)}_{k,j})=\exp\left(-(s^{(t-1)}_{k,i}-s^{(t-1)}_{k,j})^{2}/\sigma^{2}\right)$ with bandwidth $\sigma>0$, measuring how close $v_{j}$ is to $v_{i}$ in the real line spanned by $\{s^{(t-1)}_{k}\}$. As shown in the right part of Fig. 1, we then obtain an aggregated embedding for each node $v_{i}$ by a smooth weighted sum of all node embeddings:
@@
To jointly train the GNN parameters $\theta_{k}$ and the APV $\bm{a}_{k}$, we associate each node $v_{i}$ with two embeddings: $h^{(t-1)}_{k,i}$ produced by the GNN defined in Eq. 1, and $z^{(t-1)}_{k,i}$ generated via our kernel-based aggregation method as Eq. 6. We then concatenate these embeddings to form $v_{i}$’s final embedding $\pi^{(t-1)}_{k,i}=[h^{(t-1)}_{k,i}\|z^{(t-1)}_{k,i}]$, which is fed into a simple MLP classifier $\mathrm{CLF}(\cdot)$ to produce logits for the cross-entropy loss $\mathcal{L}_{k}=\frac{1}{N_{k}}CE(\mathrm{CLF}(\Gamma^{(t-1)}_{k}),Y_{k})$, where $\Gamma^{(t-1)}=[\pi^{(t-1)}_{k,i}]_{i=1}^{N_{k}}$.

### v2

@@ Abstract
Federated Learning (FL) on graph-structured data typically faces non-IID challenges, particularly in scenarios where each client holds a distinct subgraph sampled from a global graph. In this paper, we introduce Federated learning with Auxiliary projections (FedAux), a personalized subgraph FL framework that learns to align, compare, and aggregate heterogeneously distributed local models without sharing raw data or node embeddings. In FedAux, each client jointly trains (i) a local GNN and (ii) a learnable auxiliary projection vector (APV) that differentiably projects node embeddings onto a 1D space. A soft-sorting operation followed by a lightweight 1D convolution refines these embeddings in the ordered space, enabling the APV to effectively capture client-specific information. After local training, these APVs serve as compact signatures that the server uses to compute inter‑client similarities and perform similarity‑weighted parameter mixing, yielding personalized models while preserving cross‑client knowledge transfer. Moreover, we provide rigorous theoretical analysis to establish the convergence and rationality of our design. Empirical evaluations across diverse graph benchmarks demonstrate that FedAux substantially outperforms existing baselines in both accuracy and personalization performance. The code is available at https://github.com/JhuoW/FedAux.
@@ 1 Introduction
Real‑world data often manifests as relational structures, ranging from social interactions [30, 46] and financial networks [28, 45] to molecular graphs [35, 44], whose scale and privacy constraints increasingly require training to be carried out in a federated manner [9], whereby multiple clients collaboratively learn a Graph Neural Network (GNN) model without exchanging their raw data. However, applying federated learning to graph-structured data, such as social networks, faces severe challenges due to non-identically and independently distributed (non-IID) data across clients. For example, consider a federated learning scenario involving multiple regional social networking platforms, each representing a distinct subgraph of a global social network. Users within each region exhibit unique interaction patterns and distinct interests, resulting in significant heterogeneity in local graph structures and node attributes. This inherent diversity among subgraphs leads to substantial difficulties when attempting to aggregate local GNN models into a unified global model, as traditional FL algorithms [20, 14] typically assume homogeneous data distributions across clients.
To tackle the non-IID challenges inherent in subgraph federated learning, personalized FL [29] has recently emerged as a promising paradigm, which aims to provide client-specific GNN models rather than enforcing a universal global solution. Existing personalized subgraph FL approaches commonly achieve personalization by clustering clients on the server side, necessitating a reliable measure of client similarity without direct access to client-side data. In this work, we impose even stricter privacy constraints: neither raw data nor embeddings are shared, and only model learnable parameters can be exchanged. Although the server could compare clients by directly measuring similarity between their parameter matrices uploaded, the high dimensionality of these matrices makes such metrics unreliable under the curse of dimensionality [4]. Recent improvements have proposed measuring similarity by comparing communication‑level parameter gradients [35] or generating a common anchor graph on the server as a neutral testbed [3]. Although these strategies mitigate some limitations, they remain largely heuristic and do not explicitly model the heterogeneity inherent in subgraph clients (See extended discussion in Appendix˜A).
Figure 1: The overall framework of FedAux. Left: The server maintains a global GNN model together with learnable auxiliary projection vectors (APVs) that are broadcast to all clients at the start of each communication round. Middle: Clients jointly optimize the GNN and APV during local training, where the APV projects node embeddings onto a 1D $\bm{a}_{k}$-space that positions related nodes closer together. Right: After local training, clients transmit their optimized GNN parameters and personalized APVs to the server. The server computes a client similarity matrix by comparing the learned APVs, which captures the heterogeneity across subgraphs without accessing raw data. These similarities determine personalized aggregation weights. At the end of each round, the updated global parameters are broadcast for the next communication round.
@@
Contribution In this work, we propose FedAux, which employs differentiable auxiliary projections to effectively capture and exploit client-specific heterogeneity for subgraph FL. As illustrated on the left of Fig.˜1, the server stores not only a global GNN but also a learnable auxiliary projection vector (APV) that accompanies the model parameters. At the start of the first communication round, the server broadcasts the global GNN and the current APV to all clients. Each client projects its node embeddings onto the APV, which is treated as a one-dimensional latent space. A differentiable soft-sorting operator then orders the projected embeddings by similarity, after which a simulated 1D convolution aggregates the sorted embedding sequence. The aggregated representations drive a supervised loss that simultaneously refines the local GNN and the APV, so that the optimized APV preserves the relational structure of the client subgraph. Upon completing local training, clients send their updated GNN weights and personalized APVs back to the server. Since the APV reveals only the latent space that best preserves local node relationships while concealing the exact position of every node in this space, it acts as a compact privacy-preserving summary of the client subgraph. Then the server computes similarity among the returned APVs to quantify inter-client affinity and yields client-specific aggregation weights. The server uses these weights to combine the incoming parameters, producing a personalized model for each client that respects both shared knowledge and local subgraph idiosyncrasies.
@@ 2 Problem Statement: Subgraph Federated Learning
In Federated Learning (FL), multiple clients collaboratively train a global model without exchanging their raw data. In the subgraph federated learning setting, each client holds a subgraph of a larger graph. Formally, let a graph $\mathcal{G}$ be partitioned (or subsampled) into $K$ subgraphs $\{G_{1},G_{2},\cdots,G_{K}\}$ as $K$ clients, where $G_{k}=(V_{k},E_{k},X_{k},Y_{k})$. Here, $V_{k}=\{v_{k,1},\cdots,v_{k,N_{k}}\}$ is the set of nodes in the $k$-th subgraph with $N_{k}=|V_{k}|$ nodes, $E_{k}$ the set of edges among those nodes, $X_{k}\in\mathbb{R}^{N_{k}\times d}$ the node features, and $Y_{k}$ the labels relevant to the learning task. In our FL scenario, each client $G_{i}$ has access only to its local data (i.e., its subgraph structure, node features, and labels), and there is no sharing of raw data or any node embeddings between clients.
A typical GNN $f_{\theta_{k}}(G_{k})$ parameterized by $\theta_{k}$ is employed to produce node embeddings and ultimately generate predictions on the client $G_{k}$. In a standard federated learning setting such as FedAvg [20], one aims to solve the global objective: $\min_{\theta}\sum_{k=1}^{K}\alpha_{k}\mathcal{L}_{k}(\theta)$, subject to the privacy constraint that raw local data $G_{k}$ never leaves the client side. A common choice is to weight client $G_{k}$ by $\alpha_{k}=N_{k}/\bigl(\sum_{j=1}^{K}N_{j}\bigr)$ or simply $\alpha_{k}=1/K$. The iterative procedure proceeds as follows. First, the server initializes $\theta^{(0)}$. At each global communication round $t\in\{1,\cdots,T\}$, it sends $\theta^{(t-1)}$ to each client. $G_{k}$ then updates $\theta^{(t-1)}$ locally by taking a few stochastic gradient steps on $\mathcal{L}_{k}(\theta)$ to update the parameters $\theta_{k}\leftarrow\theta_{k}-\eta\nabla\mathcal{L}$, which produce $G_{k}$’s optimal local parameters $\theta_{k}^{(t)}$. After the $t$-th local training, all clients’ locally updated parameters $\{\theta_{1}^{(t)},\cdots\theta_{K}^{(t)}\}$ are sent back to the server, which aggregates them via a weighted average $\theta^{(t)}=\sum_{k=1}^{K}\alpha_{k}\theta_{k}^{(t)}$. The newly aggregated global parameters $\theta^{(t)}$ are then broadcast back to each client for the next communication round. When the process converges or reaches a designated number of rounds, the final global parameters $\theta^{(T)}$ are taken as the parameters of the GNN model on the server.
@@ 3 Methodology
In this section, we introduce the proposed Subgraph Federated Learning with Auxiliary Projections (FedAux) framework, designed to address the heterogeneity across local subgraphs in federated learning. Fig.˜1 illustrates an overview of FedAux. Our objective is twofold: 1) Each client locally encodes its subgraph into a one-dimensional space via a learnable auxiliary projection vector APV, and 2) the server then exploits these auxiliary vectors to realize personalized aggregation.
@@ 3.1 Client-Side Local Training
where $d^{\prime}$ is the output dimension of node embeddings. Given the local APV $\bm{a}_{k}^{(t-1)}$, we first normalize all node embeddings so that all embeddings are compared on a consistent scale as $\hat{h}^{(t-1)}_{k,i}=h^{(t-1)}_{k,i}/\max_{j}\|h^{(t-1)}_{k,j}\|$. Then the similarity between node $v_{k,i}$ and $\bm{a}_{k}^{(t-1)}$ is defined as $s^{(t-1)}_{k,i}=\langle\hat{h}^{(t-1)}_{k,i},\bm{a}_{k}^{(t-1)}\rangle$, where $\langle\cdot,\cdot\rangle$ denotes the inner product in $\mathbb{R}^{d^{\prime}}$. Intuitively, $s^{(t-1)}_{k,i}$ can be interpreted as the coordinate of each node $v_{k,i}$ in $\bm{a}_{k}$-space, which is a 1D line. Since $\bm{a}^{(t-1)}_{k}$ is itself learnable, the client $G_{k}$ is adaptively refining this space to capture relationships among its node embeddings more effectively.
@@
which aligns node embeddings according to their coordinates in the $\bm{a}_{k}$-space. As shown in Fig.˜2, during local training, the node embeddings and the structure of the $\bm{a}_{k}$-space are jointly optimized so that nodes with stronger relationships are positioned closer along this learned space (in $G_{k}$ comprising two triangles, nodes within the same triangle should be proximate in the $\bm{a}_{k}$-space). In other words, the objective of the local model is to adaptively reshape the APV so that the induced node sorting effectively captures the local data information.
Under the semi-supervised setting, the node sorting on APV can be adaptively refined under the guidance of the downstream task. Thinking of each $h^{(t-1)}_{k,\pi_{k}(i)}$ as a feature vector in a 1D sequence $\widetilde{\mathbf{H}}_{k}^{(t-1)}$, inspired by [18], we can apply a 1D convolution with a fixed kernel size $B$ over $\widetilde{\mathbf{H}}_{k}^{(t-1)}$ as $\mathrm{Conv1D}(\left[h^{(t-1)}_{k,\pi_{k}(1)},h^{(t-1)}_{k,\pi_{k}(2)},\cdots,h^{(t-1)}_{k,\pi_{k}(N_{k})}\right])$. More specifically, for each node $v_{\pi_{k}(i)}$ in the sorted sequence, the convolution can be written as:
@@
where $\mathbf{W}_{\tau}\in\mathbb{R}^{d^{\prime}\times d^{\prime}}$ are learnable convolution kernels for offset $\tau$, and $b$ is a bias term. Convolving around $v_{\pi_{k}(i)}$ in Eq.˜4 amounts to a proximity-based aggregation in the $\bm{a}_{k}$-space, where each embedding is updated by aggregating information from its neighbors along this learned 1D sorting. Hence, the quality of this sorting significantly impacts the aggregation effectiveness. Intuitively, nodes with stronger semantic relationships or similar labels should appear closer together in this learned sorting. We can formulate the learning objective to explicitly optimize the sorting induced by the APV $\bm{a}_{k}$, ensuring that the resultant sorting facilitates effective aggregation and improves downstream predictive accuracy. Consequently, we formulate the learning objective:
@@
$\left(\theta_{k}^{*},\bm{a}_{k}^{*},\Phi_{k}^{*}\right)=\underset{\theta_{k},a_{k},\Phi_{k}}{\operatorname{argmin}}\mathcal{L}\left(\operatorname{CLF}\left(\mathrm{Conv1D}\left(\tilde{\mathbf{H}}_{k}^{(t-1)}\right)\right),Y_{k}\right),$ (5)
@@
where $\Phi_{k}$ denotes the full set of parameters for $\mathrm{Conv1D}$ and the subsequent classifier $\operatorname{CLF}$ that maps node embeddings to final logits. Through this objective, we explicitly encourage $\bm{a}_{k}$-space to yield an optimal node sorting, enabling the convolutional operation to effectively capture and leverage localized, label-informed relationships, thereby the optimized APV $\bm{a}^{(t)}_{k}=\bm{a}^{*}_{k}$ accurately preserves and encodes the local node relationships specific to each client.
@@
However, $\bm{a}_{k}$ does not directly participate in the loss defined in Eq.˜5 in a way that enables standard backpropagation to update it. It is because the role of $\bm{a}_{k}$ is limited to generating similarity scores, which in turn determine the input order to the $\mathrm{Conv1D}$ layer. Thus, $\bm{a}_{k}$ only affects the network’s output by reordering embeddings, which is a purely indirect pathway that does not produce a gradient signal for $\bm{a}_{k}$ from the downstream loss.  [18] attempted to mitigate this by multiplying each node embedding by its similarity score and then sorting. While this modification integrates $\bm{a}_{k}$ into the learning pipeline directly, the hard discrete sort persists, causing the gradient signal that could refine $\bm{a}_{k}$ to be still routed through a non-smooth transformation. Hence $\bm{a}_{k}$ still cannot be fully optimized to reorder embeddings based on loss feedback, leaving the core issue unresolved.
@@
To eliminate the hard-sorting bottleneck, we propose a continuous aggregation scheme over the $\bm{a}_{k}$-space. Rather than ranking or discretizing these similarity scores, for each node $v_{i}$, we define a continuous kernel $\kappa(s^{(t-1)}_{k,i},s^{(t-1)}_{k,j})$, which could be a simple Gaussian-like function $\mathcal{K}_{ij}=\kappa(s^{(t-1)}_{k,i},s^{(t-1)}_{k,j})=\exp\left(-(s^{(t-1)}_{k,i}-s^{(t-1)}_{k,j})^{2}/\sigma^{2}\right)$ with bandwidth $\sigma>0$, measuring how close $v_{j}$ is to $v_{i}$ in the real line spanned by $\{s^{(t-1)}_{k}\}$. As shown in the right part of Fig.˜1, we then obtain an aggregated embedding for each node $v_{i}$ by a smooth weighted sum of all node embeddings:
@@
To jointly train the GNN parameters $\theta_{k}$ and the APV $\bm{a}_{k}$, we associate each node $v_{i}$ with two embeddings: $h^{(t-1)}_{k,i}$ produced by the GNN defined in Eq.˜1, and $z^{(t-1)}_{k,i}$ generated via our kernel-based aggregation method as Eq.˜6. We then concatenate these embeddings to form $v_{i}$’s final embedding $r^{(t-1)}_{k,i}=[h^{(t-1)}_{k,i}\|z^{(t-1)}_{k,i}$], which is fed into a simple MLP classifier $\mathrm{CLF}(\cdot)$ to produce logits for the cross-entropy loss $\mathcal{L}_{k}=\frac{1}{N_{k}}CE(\mathrm{CLF}(\Gamma^{(t-1)}_{k}),Y_{k})$, where $\Gamma^{(t-1)}=[r^{(t-1)}_{k,i}]_{i=1}^{N_{k}}$.